Given an x-mean of 13.8 and a y-mean of 3.1, calculate the crossproduct of x=15 and y=6.
Answer #1

solution Crossproduct of x and y is c.p = (x-7)(y ý given x = 15 y = 6 Mean n = = 13.8 Mean 3.1 y اعه C.P= (15-13.8) (6-3-1)any query, please ask